what is coulombs inverse square law

Grade:12

1 Answers

Askiitian.Expert Rajat
14 years ago

HI,

Coulomb's law, sometimes called the Coulomb law, is an equation describing the electrostatic force between electric charges. It was developed in the 1780s by French physicist Charles Augustin de Coulomb and was essential to the development of the theory of electromagnetism. Coulomb's law may be stated in scalar form as follows:

    The magnitude of the electrostatic force between two point electric charges is directly proportional to the product of the magnitudes of each of the charges and inversely proportional to the square of the total distance between the two charges.

The scalar form of Coulomb's law will only describe the magnitude of the electrostatic force between two electric charges. If direction is required, then the vector form is required as well. The magnitude of the electrostatic force (F) on a charge (q1) due to the presence of a second charge (q2), is given by

F = k_\mathrm{e} \frac{q_1q_2}{r^2},


where r is the distance between the two charges and ke a proportionality constant. A positive force implies a repulsive interaction, while a negative force implies an attractive interaction.

The proportionality constant ke, called Coulomb's constant is related to the properties of space and can be calculated exactly:

\begin{align} k_{\mathrm{e}} &= \frac{1}{4\pi\varepsilon_0} = \frac{\mu_0\ {c_0}^2}{4 \pi}= \frac{{c_0}^2}{10^7}=\\ &= 8.987\ 551\ 787\ 368\ 176\ 4 \times 10^9 \ \mathrm{N  \cdot m^2 \cdot C^{-2}}. \\ \end{align}